# Graph Laplacian-based Bayesian Multi-fidelity Modeling

### v1

@@ 3.2.1 Nyström-QR approximation
$\bm{\Theta}$ $:=\bm{P}_{M}^{T}\bm{P}_{M}+\sigma^{2}\omega(1+\tau)^{\beta}\bm{I}_{N}\in\mathbb{R}^{N\times N},$
$\bm{\Xi}$ $:=\sigma^{2}\omega\left((1+\tau)^{\beta}\bm{I}_{K}-((1+\tau)\bm{I}_{K}-\bm{\Sigma})^{\beta}\right)\in\mathbb{R}^{K\times K},$
it follows that Equation 2.11 can be approximately solved by solving
$\left(\bm{\Theta}-\tilde{\bm{U}}\bm{\Xi}\tilde{\bm{U}}^{T}\right)\bm{\Phi}^{*}=\bm{P}_{M}^{T}\hat{\bm{\Phi}},$ (3.5)
@@
$\begin{pmatrix}\bm{\Theta}&\tilde{\bm{U}}\\ \tilde{\bm{U}}^{T}&\bm{\Xi}^{-1}\end{pmatrix}\begin{pmatrix}\bm{\Phi}^{*}\\ \bm{B}\end{pmatrix}$ $=\begin{pmatrix}\bm{P}_{M}^{T}\hat{\bm{\Phi}}\\ \bm{0}\end{pmatrix}$ (3.6)
@@
$\begin{pmatrix}\bm{\Theta}&\tilde{\bm{U}}\\ \bm{\Xi}\tilde{\bm{U}}^{T}&\bm{I}_{K}\end{pmatrix}\begin{pmatrix}\bm{\Phi}^{*}\\ \bm{B}\end{pmatrix}$ $=\begin{pmatrix}\bm{P}_{M}^{T}\hat{\bm{\Phi}}\\ \bm{0}\end{pmatrix},$ (3.7)
@@
$\bm{\Theta}\bm{\Phi}^{*}+\tilde{\bm{U}}\bm{B}$ $=\bm{P}_{M}^{T}\hat{\bm{\Phi}},$ $\bm{B}=-\bm{\Xi}\tilde{\bm{U}}^{T}\bm{\Phi}^{*}.$
@@
$\begin{pmatrix}\bm{\Theta}&\tilde{\bm{U}}\\ \tilde{\bm{U}}^{T}&\bm{\Xi}^{-1}\end{pmatrix}$ and $\begin{pmatrix}\bm{\Theta}&\tilde{\bm{U}}\\ \bm{\Xi}\tilde{\bm{U}}^{T}&\bm{I}_{K}\end{pmatrix}$
are extremely sparse, with at most $2NK+N+K$ non-zero entries, Equation 3.6 and Equation 3.7 can then be efficiently solved by sparse linear solvers, e.g. GMRES, see Saad [57] for details. Since Equation 3.6 is symmetric as well as sparse, it can be solved more efficiently than Equation 3.7, e.g. via MINRES [58], so long as the $\bm{\Xi}^{-1}$ term causes no issues.
@@
i.e., $\sigma_{i}\approx 0$ or $\beta$ is an even integer and $\sigma_{i}\approx 2(1+\tau)$. Since the $\sigma_{i}$ are approximate eigenvalues for $\bm{I}_{N}-\bm{L}$, they will approximately lie in $[-1,1]$, so for $\tau\geq 0$ this latter case will not arise so long as the approximation is sufficiently good, and furthermore $\sigma_{i}\leq 1+\tau$ is desirable anyway so that $\bm{\Xi}$ is defined for all $\beta\geq 0$. The case of $\sigma_{i}\approx 0$ can be rectified by checking for this in advance and then redefining $\tilde{\bm{U}}$ and $\bm{\Sigma}$ with those columns of $\tilde{\bm{U}}$ and rows/columns of $\bm{\Sigma}$ removed.
@@ 3.2.3 Computing the covariance matrix
$\bm{C}=\left(\frac{1}{\sigma^{2}}\bm{P}_{M}^{T}\bm{P}_{M}+\omega(\bm{L}+\tau\bm{I}_{N})^{\beta}\right)^{-1}.$
@@
$\begin{split}\bm{C}&\approx\sigma^{2}\left(\bm{\Theta}-\tilde{\bm{U}}\bm{\Xi}\tilde{\bm{U}}^{T}\right)^{-1}\\ &=\sigma^{2}\left(\bm{\Theta}^{-1}+\bm{\Theta}^{-1}\tilde{\bm{U}}(\bm{\Xi}^{-1}-\tilde{\bm{U}}^{T}\bm{\Theta}^{-1}\tilde{\bm{U}})^{-1}\tilde{\bm{U}}^{T}\bm{\Theta}^{-1}\right),\end{split}$ (3.8)
with the latter equality following from the Woodbury identity [59]. This latter term requires $\mathcal{O}(N^{2}K)$ time to compute and $\mathcal{O}(N^{2})$ space to store, so may be intractable. However, if all that are desired are matrix-vector products with $\bm{C}$, then these can be computed by precomputing
$\left(\bm{\Xi}^{-1}-\tilde{\bm{U}}^{T}\bm{\Theta}^{-1}\tilde{\bm{U}}\right)^{-1}$
which requires $\mathcal{O}(NK^{2})$ time, and then for any vector $\bm{v}\in\mathbb{R}^{N}$, $\bm{C}\bm{v}$ can be approximated via (3.8) in $\mathcal{O}(NK)$ time.
@@
The expression in Equation 3.8 provides another method for solving Equation 3.5.
@@ 3.2.4 Computational complexities
Quantity Complexity in time Complexity in space
$\bm{W}(:,X)$ $\mathcal{O}(NK)$ $\mathcal{O}(NK)$
$\bm{W}(X,X)^{\dagger}$ $\mathcal{O}(K^{3})$ $\mathcal{O}(K^{2})$
$\hat{\bm{D}}$ $\mathcal{O}(NK)$ $\mathcal{O}(N)$
$\bm{Q}$ $\mathcal{O}(NK^{2})$ $\mathcal{O}(NK)$
$\bm{R}$ $\mathcal{O}(NK^{2})$ $\mathcal{O}(K^{2})$
$\bm{\Gamma,\Sigma}$ $\mathcal{O}(K^{3})$ $\mathcal{O}(K^{2})$
$\tilde{\bm{U}}$ $\mathcal{O}(NK^{2})$ $\mathcal{O}(NK)$
$\bm{\Theta}$ $\mathcal{O}(N)$ $\mathcal{O}(N)$
$\bm{\Xi}$ $\mathcal{O}(K)$ $\mathcal{O}(K)$
$\bm{\Phi}^{*}$ ($m$-step GMRES on Equation 3.7) $\mathcal{O}((m^{2}N+mNK)D)$ $\mathcal{O}(mN+m^{2}+ND)$
$\bm{\Phi}^{*}$ ($m$-step MINRES on Equation 3.6) $\mathcal{O}(mNKD)$ $\mathcal{O}(m^{2}+ND)$
$\bm{\Phi}^{*}$ (via Woodbury Equation 3.8) $\mathcal{O}(NK^{2}+NKD)$ $\mathcal{O}(ND)$
$\bm{C}$ $\mathcal{O}(N^{2}K)$ $\mathcal{O}(N^{2})$
$\bm{C}\bm{v}$ $\mathcal{O}(NK^{2})$ (once)${}+\mathcal{O}(NK)$ (per $\bm{v}$) $\mathcal{O}(N+K^{2})$
@@ 4 Convergent regularization
We here prove a convergent regularization result. This result shows that as the error in the high-fidelity data tends to zero, one can correspondingly tune down the regularization strength $\omega$ such that the corresponding MAP displacements converge to a displacement which agrees with the true displacement on the high-fidelity data. This gives a guarantee against over-regularization. Our proof technique in this section will be a special case of the technique employed in Pöschl [60].
We begin by noting that for a fixed $\omega$, which denotes the strength of the prior, finding the MAP estimate $\bm{\Phi}^{*}$ Equation 2.10 is equivalent to solving the optimization problem
$\bm{\Phi}^{*}\in\operatorname*{argmin}_{\bm{\Theta}\in\mathbb{R}^{N\times D}}\frac{1}{2\sigma^{2}}\|\bm{P}_{M}\mathbf{\Theta}-\hat{\bm{\Phi}}\|^{2}_{F}+\frac{1}{2}\omega\langle\mathbf{\Theta},(\bm{L}+\tau\bm{I}_{N})^{\beta}\mathbf{\Theta}\rangle_{F}.$
@@
$\bm{\Phi}_{n}^{*}\in\operatorname*{argmin}_{\bm{\Theta}\in\mathbb{R}^{N\times D}}\frac{1}{2}\|\bm{P}_{M}\bm{\Theta}-\hat{\bm{\Phi}}_{n}\|^{2}_{F}+\omega_{n}\underbrace{\langle\mathbf{\Theta},(\bm{L}+\tau\bm{I}_{N})^{\beta}\mathbf{\Theta}\rangle_{F}}_{=:\mathcal{R}(\bm{\Theta})}.$
Thus, $\bm{\Phi}_{n}^{*}$ is the MAP estimator for the true displacements $\bm{\Phi}$, given the $n^{\text{th}}$ high-fidelity data $\{\bm{\hat{u}}^{(i)}_{n}\}_{i=1}^{M}$ and setting $\omega=\omega_{n}$. We will show that if the error in the high-fidelity data vanishes and the $\omega_{n}$ are chosen with appropriate decay (i.e., tending to zero but more slowly than the squared error in the high-fidelity data) then the $\bm{\Phi}_{n}^{*}$ will converge to a displacement matrix which agrees with the true displacements at $1$ to $M$, and minimises $\mathcal{R}$ among the displacement matrices satisfying that constraint.
@@
  • $\hat{\bm{\Phi}}_{n}=\bm{P}_{M}\bm{\Phi}+\bm{E}_{n}$ where $[\bm{E}_{n}]_{ij}=(\bm{\hat{u}}_{n})_{j}^{(i)}-\bm{u}^{(i)}_{j}$ for $i\in\{1,...,M\}$ and $j\in\{1,...,D\}$, i.e. $\bm{E}_{n}$ describes the error in the $n^{\text{th}}$ realization of the high-fidelity data.
@@
[…]
@@
$\operatorname*{argmin}_{\bm{\Theta}\in\mathbb{R}^{N\times D}}\mathcal{R}(\bm{\Theta})\qquad\text{ s.t. }\qquad\bm{P}_{M}\bm{\Theta}=\bm{P}_{M}\bm{\Phi}.$ (4.1)
@@
Since $\mathcal{R}$ is strictly convex, continuous, coercive, and bounded below it follows that Equation 4.1 has a unique solution $\bm{\Phi}_{\infty}^{*}$. By 4.1, it will suffice to show that every subsequence of $\bm{\Phi}_{n}^{*}$ has a convergent subsubsequence, and that the limit of this subsubsequence solves Equation 4.1 and is thus equal to $\bm{\Phi}_{\infty}^{*}$.
@@
$0\leq\omega_{n}\mathcal{R}(\bm{\Phi}_{n}^{*})+\frac{1}{2}\|\bm{P}_{M}\bm{\Phi}_{n}^{*}-\bm{\hat{\Phi}}_{n}\|_{F}^{2}\leq\omega_{n}\mathcal{R}(\bm{\Phi})+\frac{1}{2}\|\bm{P}_{M}\bm{\Phi}-\bm{\hat{\Phi}}_{n}\|_{F}^{2}\to 0$ (4.2)
@@
$0\leq\mathcal{R}(\bm{\Phi}_{n}^{*})\leq\mathcal{R}(\bm{\Phi}_{n}^{*})+\frac{1}{2\omega_{n}}\|\bm{P}_{M}\bm{\Phi}_{n}^{*}-\bm{\hat{\Phi}}_{n}\|_{F}^{2}\leq\mathcal{R}(\bm{\Phi})+\frac{\delta_{n}^{2}}{2\omega_{n}}\to\mathcal{R}(\bm{\Phi}),$
@@
Therefore, every subsequence of $\bm{\Phi}_{n}^{*}$ has a convergent subsubsequence. Abusing notation, we will denote an arbitrary such subsubsequence by $\bm{\Phi}_{n}^{*}$ (and likewise for the corresponding subsubsequences of $\omega_{n}$, $\bm{\hat{\Phi}}_{n}$, and $\delta_{n}$), and denote its limit by $\bm{\tilde{\Phi}}$. It remains to show that $\bm{\tilde{\Phi}}$ solves Equation 4.1.
We first show that $\bm{P}_{M}\bm{\tilde{\Phi}}=\bm{P}_{M}\bm{\Phi}$. Note that $\bm{P}_{M}\bm{\tilde{\Phi}}=\lim_{n\to\infty}\bm{P}_{M}\bm{\Phi}_{n}^{*}$, that by Equation 4.2 and the nonnegativity of $\mathcal{R}$ we have that $\|\bm{P}_{M}\bm{\Phi}_{n}^{*}-\bm{\hat{\Phi}}_{n}\|_{F}\to 0$, and that by (i) $\bm{\hat{\Phi}}_{n}\to\bm{P}_{M}\bm{\Phi}$. The claim follows.
@@
$\omega_{n}\mathcal{R}(\bm{\Phi}_{n}^{*})$ $\leq\omega_{n}\mathcal{R}(\bm{\Phi}_{n}^{*})+\frac{1}{2}\|\bm{P}_{M}\bm{\Phi}_{n}^{*}-\bm{\hat{\Phi}}_{n}\|_{F}^{2}$
$\leq\omega_{n}\mathcal{R}(\bm{\Theta})+\frac{1}{2}\|\bm{P}_{M}\bm{\Theta}-\bm{\hat{\Phi}_{n}}\|_{F}^{2}$
$=\omega_{n}\mathcal{R}(\bm{\Theta})+\frac{1}{2}\delta_{n}^{2}$
@@
$\mathcal{R}(\bm{\tilde{\Phi}})=\lim_{n\to\infty}\mathcal{R}(\bm{\Phi}_{n}^{*})\leq\lim_{n\to\infty}\mathcal{R}(\bm{\Theta})+\frac{\delta_{n}^{2}}{2\omega_{n}}=\mathcal{R}(\bm{\Theta}).$

### v2

@@ 3.2.2 Solving (20)
$\bm{\Theta}$ $:=\bm{P}_{M}^{T}\bm{P}_{M}+\sigma^{2}\omega(1+\tau)^{\beta}\bm{I}_{N}\in\mathbb{R}^{N\times N},$ (43)
$\bm{\Xi}$ $:=\sigma^{2}\omega\left((1+\tau)^{\beta}\bm{I}_{K}-((1+\tau)\bm{I}_{K}-\bm{\Sigma})^{\beta}\right)\in\mathbb{R}^{K\times K},$ (44)
it follows that Equation 20 can be approximately solved by solving
$\left(\bm{\Theta}-\tilde{\bm{U}}\bm{\Xi}\tilde{\bm{U}}^{T}\right)\bm{\Phi}^{*}=\bm{P}_{M}^{T}\hat{\bm{\Phi}}$, (45)
@@
$\begin{pmatrix}\bm{\Theta}&\tilde{\bm{U}}\\ \tilde{\bm{U}}^{T}&\bm{\Xi}^{-1}\end{pmatrix}\begin{pmatrix}\bm{\Phi}^{*}\\ \bm{B}\end{pmatrix}$ $=\begin{pmatrix}\bm{P}_{M}^{T}\hat{\bm{\Phi}}\\ \bm{0}\end{pmatrix}$ (46)
@@
$\begin{pmatrix}\bm{\Theta}&\tilde{\bm{U}}\\ \bm{\Xi}\tilde{\bm{U}}^{T}&\bm{I}_{K}\end{pmatrix}\begin{pmatrix}\bm{\Phi}^{*}\\ \bm{B}\end{pmatrix}$ $=\begin{pmatrix}\bm{P}_{M}^{T}\hat{\bm{\Phi}}\\ \bm{0}\end{pmatrix}$, (47)
@@
$\bm{\Theta}\bm{\Phi}^{*}+\tilde{\bm{U}}\bm{B}$ $=\bm{P}_{M}^{T}\hat{\bm{\Phi}}$, $\bm{B}=-\bm{\Xi}\tilde{\bm{U}}^{T}\bm{\Phi}^{*}.$ (48)
@@
$\begin{pmatrix}\bm{\Theta}&\tilde{\bm{U}}\\ \tilde{\bm{U}}^{T}&\bm{\Xi}^{-1}\end{pmatrix}$ and $\begin{pmatrix}\bm{\Theta}&\tilde{\bm{U}}\\ \bm{\Xi}\tilde{\bm{U}}^{T}&\bm{I}_{K}\end{pmatrix}$ (49)
are extremely sparse, with at most $2NK+N+K$ non-zero entries, Equation 46 and Equation 47 can then be efficiently solved by sparse linear solvers, e.g. GMRES, see Saad [Saad2003] for details. Since Equation 46 is symmetric as well as sparse, it can be solved more efficiently than Equation 47, e.g. via MINRES [MINRES], so long as the $\bm{\Xi}^{-1}$ term causes no issues.
@@
i.e., $\sigma_{i}\approx 0$ or $\beta$ is an even integer and $\sigma_{i}\approx 2(1+\tau)$. Since the $\sigma_{i}$ are approximate eigenvalues for $\bm{I}_{N}-\bm{L}$, they will approximately lie in [1,1][-1,1], so for $\tau\geq 0$ this latter case will not arise so long as the approximation is sufficiently good, and furthermore $\sigma_{i}\leq 1+\tau$ is desirable anyway so that $\bm{\Xi}$ is defined for all $\beta\geq 0$. The case of $\sigma_{i}\approx 0$ can be rectified by checking for this in advance and then redefining $\tilde{\bm{U}}$ and $\bm{\Sigma}$ with those columns of $\tilde{\bm{U}}$ and rows/columns of $\bm{\Sigma}$ removed.
@@ 3.2.3 Computing the covariance matrix
$\bm{C}=\left(\frac{1}{\sigma^{2}}\bm{P}_{M}^{T}\bm{P}_{M}+\omega(\bm{L}+\tau\bm{I}_{N})^{\beta}\right)^{-1}.$ (50)
@@
$\begin{split}\bm{C}&\approx\sigma^{2}\left(\bm{\Theta}-\tilde{\bm{U}}\bm{\Xi}\tilde{\bm{U}}^{T}\right)^{-1}\\ &=\sigma^{2}\left(\bm{\Theta}^{-1}+\bm{\Theta}^{-1}\tilde{\bm{U}}(\bm{\Xi}^{-1}-\tilde{\bm{U}}^{T}\bm{\Theta}^{-1}\tilde{\bm{U}})^{-1}\tilde{\bm{U}}^{T}\bm{\Theta}^{-1}\right),\end{split}$ (51)
with the latter equality following from the Woodbury identity [Woodbury50]. This latter term requires $\mathcal{O}(N^{2}K)$ time to compute and $\mathcal{O}(N^{2})$ space to store, so may be intractable. However, if all that are desired are matrix-vector products with $\bm{C}$, then these can be computed by precomputing
$\left(\bm{\Xi}^{-1}-\tilde{\bm{U}}^{T}\bm{\Theta}^{-1}\tilde{\bm{U}}\right)^{-1}$ (52)
which requires $\mathcal{O}(NK^{2})$ time, and then for any vector $\bm{v}\in\mathbb{R}^{N}$, $\bm{C}\bm{v}$ can be approximated via (51) in $\mathcal{O}(NK)$ time.
@@
The expression in Equation 51 provides another method for solving Equation 45.
@@ 3.2.4 Computational complexities
[…]
@@ 4 Convergent regularization
We here prove a convergent regularization result. This result shows that as the error in the high-fidelity data tends to zero, one can correspondingly tune down the regularization strength $\omega$ such that the corresponding MAP displacements converge to a displacement which agrees with the true displacement on the high-fidelity data. This gives a guarantee against over-regularization. Our proof technique in this section will be a special case of the technique employed in Pöschl [poschl2008tikhonov].
We begin by noting that for a fixed $\omega$, which denotes the strength of the prior, finding the MAP estimate $\bm{\Phi}^{*}$ Equation 19 is equivalent to solving the optimization problem
$\bm{\Phi}^{*}\in\operatorname*{argmin}_{\bm{\Theta}\in\mathbb{R}^{N\times D}}\frac{1}{2\sigma^{2}}\|\bm{P}_{M}\mathbf{\Theta}-\hat{\bm{\Phi}}\|^{2}_{F}+\frac{1}{2}\omega\langle\mathbf{\Theta},(\bm{L}+\tau\bm{I}_{N})^{\beta}\mathbf{\Theta}\rangle_{F}.$ (53)
@@
$\bm{\Phi}_{n}^{*}\in\operatorname*{argmin}_{\bm{\Theta}\in\mathbb{R}^{N\times D}}\frac{1}{2}\|\bm{P}_{M}\bm{\Theta}-\hat{\bm{\Phi}}_{n}\|^{2}_{F}+\omega_{n}\underbrace{\langle\mathbf{\Theta},(\bm{L}+\tau\bm{I}_{N})^{\beta}\mathbf{\Theta}\rangle_{F}}_{=:\mathcal{R}(\bm{\Theta})}.$ (55)
Thus, $\bm{\Phi}_{n}^{*}$ is the MAP estimator for the true displacements $\bm{\Phi}$, given the $n^{\text{th}}$ high-fidelity data $\{\bm{\hat{u}}^{(i)}_{n}\}_{i=1}^{M}$ and setting $\omega=\omega_{n}$. We will show that if the error in the high-fidelity data vanishes and the $\omega_{n}$ are chosen with appropriate decay (i.e., tending to zero but more slowly than the squared error in the high-fidelity data) then the $\bm{\Phi}_{n}^{*}$ will converge to a displacement matrix which agrees with the true displacements at the locations $i\in\{1,...,M\}$, and minimises $\mathcal{R}$ among the displacement matrices satisfying that constraint.
@@
  • 2.
    $\hat{\bm{\Phi}}_{n}=\bm{P}_{M}\bm{\Phi}+\bm{E}_{n}$ where $\bm{E}_{n}]_{ij}=(\bm{\hat{u}}_{n})_{j}^{(i)}-\bm{u}^{(i)}_{j}$ for $i\in\{1,...,M\}$ and $j\in\{1,...,D\}$, i.e. $\bm{E}_{n}$ describes the error in the $n^{\text{th}}$ realization of the high-fidelity data.
@@
$\text{(i) }\delta_{n}\downarrow 0,$ $\text{(ii) }\omega_{n}\downarrow 0,$ and $\text{(iii) }\frac{\delta_{n}^{2}}{\omega_{n}}\to 0.$ (56)
@@
$\operatorname*{argmin}_{\bm{\Theta}\in\mathbb{R}^{N\times D}}\mathcal{R}(\bm{\Theta})\qquad\text{ s.t. }\qquad\bm{P}_{M}\bm{\Theta}=\bm{P}_{M}\bm{\Phi}.$ (57)
@@
Since $\mathcal{R}$ is strictly convex, continuous, coercive, and bounded below it follows that Equation 57 has a unique solution $\bm{\Phi}_{\infty}^{*}$. By 4.1, it will suffice to show that every subsequence of $\bm{\Phi}_{n}^{*}$ has a convergent subsubsequence, and that the limit of this subsubsequence solves Equation 57 and is thus equal to $\bm{\Phi}_{\infty}^{*}$.
@@
$0\leq\omega_{n}\mathcal{R}(\bm{\Phi}_{n}^{*})+\frac{1}{2}\|\bm{P}_{M}\bm{\Phi}_{n}^{*}-\bm{\hat{\Phi}}_{n}\|_{F}^{2}\leq\omega_{n}\mathcal{R}(\bm{\Phi})+\frac{1}{2}\|\bm{P}_{M}\bm{\Phi}-\bm{\hat{\Phi}}_{n}\|_{F}^{2}\to 0.$ (58)
@@
$0\leq\mathcal{R}(\bm{\Phi}_{n}^{*})\leq\mathcal{R}(\bm{\Phi}_{n}^{*})+\frac{1}{2\omega_{n}}\|\bm{P}_{M}\bm{\Phi}_{n}^{*}-\bm{\hat{\Phi}}_{n}\|_{F}^{2}\leq\mathcal{R}(\bm{\Phi})+\frac{\delta_{n}^{2}}{2\omega_{n}}\to\mathcal{R}(\bm{\Phi}),$ (59)
@@
Therefore, every subsequence of $\bm{\Phi}_{n}^{*}$ has a convergent subsubsequence. Abusing notation, we will denote an arbitrary such subsubsequence by $\bm{\Phi}_{n}^{*}$ (and likewise for the corresponding subsubsequences of $\omega_{n}$, $\bm{\hat{\Phi}}_{n}$, and $\delta_{n}$), and denote its limit by $\bm{\tilde{\Phi}}$. It remains to show that $\bm{\tilde{\Phi}}$ solves Equation 57.
We first show that $\bm{P}_{M}\bm{\tilde{\Phi}}=\bm{P}_{M}\bm{\Phi}$. Note that $\bm{P}_{M}\bm{\tilde{\Phi}}=\lim_{n\to\infty}\bm{P}_{M}\bm{\Phi}_{n}^{*}$, that by Equation 58 and the nonnegativity of $\mathcal{R}$ we have that $\|\bm{P}_{M}\bm{\Phi}_{n}^{*}-\bm{\hat{\Phi}}_{n}\|_{F}\to 0$, and that by (i) $\bm{\hat{\Phi}}_{n}\to\bm{P}_{M}\bm{\Phi}$. The claim follows.
@@
$\omega_{n}\mathcal{R}(\bm{\Phi}_{n}^{*})$ $\leq\omega_{n}\mathcal{R}(\bm{\Phi}_{n}^{*})+\frac{1}{2}\|\bm{P}_{M}\bm{\Phi}_{n}^{*}-\bm{\hat{\Phi}}_{n}\|_{F}^{2}$ (60)
$\leq\omega_{n}\mathcal{R}(\bm{\Theta})+\frac{1}{2}\|\bm{P}_{M}\bm{\Theta}-\bm{\hat{\Phi}_{n}}\|_{F}^{2}$ (61)
$=\omega_{n}\mathcal{R}(\bm{\Theta})+\frac{1}{2}\delta_{n}^{2}$ (62)
@@
$\mathcal{R}(\bm{\tilde{\Phi}})=\lim_{n\to\infty}\mathcal{R}(\bm{\Phi}_{n}^{*})\leq\lim_{n\to\infty}\mathcal{R}(\bm{\Theta})+\frac{\delta_{n}^{2}}{2\omega_{n}}=\mathcal{R}(\bm{\Theta}).$ (63)